# Orbital optical activity in noncentrosymmetric metals and superconductors

### v1

@@ VI Conclusion and discussion
In this paper, we have discussed the optical activity originating from the orbital magnetic moment in metals and superconductors using a Green’s function formalism. We have shown that the no-go theorem in the normal state, stating that a singular contribution to the optical activity is absent, is valid except for the chiral magnetic effect and that this theorem is broken for the superconducting state in Secs. II and III. Then, this singularity contributes to the $\delta$ function in $\mathrm{Im}\sigma_{\mu\nu\lambda}$ in the DC-limit and the anomalous $\Omega^{-1}$-divergence in $\mathrm{Re}\sigma_{\mu\nu\lambda}$ for superconductors. In the superconducting state, the coefficient of the singular part is confirmed to be exactly the superconducting Edelstein effect (SEE), which implies the validity of the missing area measurement for the SEE. An identical discussion was recently presented for the spin-induced optical activity. In this paper, it has been shown that the same discussion holds for the orbital-induced optical activity. In Sec. IV, we have shown calculations using a model for monolayer TMDs. We have used an effective two-band model to elucidate the typical behavior of the optical activity. We have revealed that the low-frequency spectrum is described by a Drude-like model and that the interband transition shows an atypical chemical potential dependence originating from the interband Fermi surface term, unique to the orbital part in metals. In addition, we have evaluated the Kerr spectrum. In Sec. V, we have calculated the spectrum of the optical activity in the superconducting state, and we have demonstrated the existence of the missing area and the anomalous divergence of the optical activity at low frequencies. In addition, we have discussed the modification of the Kerr effect by the superconducting state.
@@
Finally, we discuss the possibility of an experimental realization separated into the interband regime and the intraband regime. As discussed in Sec. IV, the interband peak is located at about a few $\mathrm{eV}$. The Kerr rotation and ellipticity for the reflected light reach a few µrad, which makes them detectable with the state-of-the-art instruments, having a resolution of 10 nrad [70, 71]. On the other hand, the resolution in the terahertz regime, corresponding to the gap scale of most noncentrosymmetric superconductors, has recently been reported to reach 20 µrad and is expected to go down to 2 µrad [72]. The Kerr effect of TBG is close to this technical observable limit. Thus, unless we can find other two-dimensional materials with a larger Edelstein coefficient $\alpha$, the experimental observation is currently challenging in this regime.
@@
Finally, we shortly discuss the three-dimensional materials. In three-dimensional systems, the Faraday effect is expected. As light cannot easily penetrate superconductors, transmission can only occur in thin films [73]. The Faraday effect originates from the difference of the refractive index for left and right circular polarized lights $\Delta n$. It is given by $\Phi_{\mathrm{F}}(=\theta_{\mathrm{F}}+i\varepsilon_{\mathrm{F}})=-\Delta nk\zeta/2$, where $\theta_{\mathrm{F}}$ and $\varepsilon_{\mathrm{F}}$ are the rotational angle and the ellipticity for the Faraday effect, $k$ is the wavenumber of the light, and $\zeta$ is the thickness of materials. For the natural optical activity, the refractive index difference is given by $\Delta n=\mu_{0}c\chi$ [74]; thus, we can directly obtain the spectrum of the optical activity $\chi$ from the Faraday spectrum. In the case of noncentrosymmetric cubics (point group: O and T), the optical activity in the THz regime is $\chi\sim(e\mu_{\mathrm{B}}\tau k_{F}^{2}/\hbar)(\gamma k_{F}/E_{F})$ with an antisymmetric SOC term $\gamma\bm{k}\cdot\bm{\sigma}$. Here, $\mu_{\mathrm{B}}$ is the Bohr magneton, $\tau$ is the relaxation time, and $k_{F}$ and $E_{F}$ are the Fermi wavenumber and the Fermi energy. Candidates for superconductors with this SOC include $\mathrm{Li_{2}Pd_{3}B}$, $\mathrm{Li_{2}Pt_{3}B}$, and $\mathrm{Mo_{2}Al_{3}C}$ [75]. Then, the Faraday effect reads
$\Phi_{\mathrm{F}}\sim-\frac{1}{2}\frac{e\mu_{\mathrm{B}}\mu_{0}k_{F}}{\hbar}(\Omega\tau)(k_{F}\zeta)\frac{\gamma k_{F}}{E_{F}},$ (21)
@@
and the typical order of magnitude is estimated to be 4.8 µrad for $k_{F}=1~{}\mathrm{\AA}^{-1}$, $\tau=10~{}\mathrm{fs}$, and $\gamma k_{F}/E_{F}=10^{-2}$ at $\Omega=1~{}\mathrm{THz}$ using the thickness $\zeta=500~{}\mathrm{\AA}$.
@@ Appendix A No-go theorem
This equation is purely imaginary and antisymmetric under the interchange between $\mu$ and $\nu$. However, there is one difference to the case of the normal state. The velocity operator appearing by differentiating the Green’s function $\hat{v}^{\mu}_{\bm{k}}$ is different from the velocity operator $\tilde{v}^{\mu}_{\bm{k}}$. Thus, many components of the singular part can be finite.
Appendix B Optical Kerr effect
@@
Here, we ignore the square of $\chi$ due to its smallness. This ratio gives the Kerr rotation $\theta$ and ellipticity $\varepsilon$ as
@@
The Kerr rotation and ellipticity of the transmitted light are obtained by simply replacing $r$ with $t$.

### v2

@@ VI.1 Conclusion
In this paper, we have discussed the optical activity originating from the orbital magnetic moment in metals and superconductors using a Green’s function formalism. We have reconfirmed that the no-go theorem in the normal state, stating that a singular contribution to the optical activity is absent, is valid except for the chiral magnetic effect and, we have shown that this theorem is broken for the superconducting state in Secs. II and III. Then, this singularity contributes to the $\delta$ function in $\mathrm{Im}\sigma_{\mu\nu\lambda}$ in the DC-limit and the anomalous $\Omega^{-1}$-divergence in $\mathrm{Re}\sigma_{\mu\nu\lambda}$ for superconductors. In the superconducting state, the coefficient of the singular part is confirmed to be exactly the superconducting Edelstein effect (SEE), which implies the validity of the missing area measurement for the SEE. An identical discussion was recently presented for the spin-induced optical activity. In this paper, it has been shown that the same discussion holds for the orbital-induced optical activity. In Sec. IV, we have shown calculations using a model for monolayer TMDs in the normal state. We have used an effective two-band model to elucidate the typical behavior of the optical activity. We have revealed that the low-frequency spectrum is described by a Drude-like model and that the interband transition shows an atypical chemical potential dependence originating from the interband Fermi surface term, unique to the orbital part in metals. We have obtained a sum rule related to the spectrum of the interband Fermi surface term. It shows that the spectrum area is determined by the orbital magnetic moment dipole. In addition, we have evaluated the spectrum of the optical rotation. In Sec. V, we have calculated the spectrum of the optical activity in the superconducting state, and we have demonstrated the existence of the missing area and the anomalous divergence of the optical activity at low frequencies. In addition, we have discussed the modification of the optical rotation by the superconducting state.
VI.2 Discussion of possible future experiments and directions
@@
Finally, we discuss the possibility of an experimental realization separated into the interband regime and the intraband regime. As discussed in Sec. IV, the interband peak is located at about a few $\mathrm{eV}$. The rotational angle and ellipticity for the reflected light reach a few µrad, which makes them detectable with the state-of-the-art instruments, having a resolution of 10 nrad [70, 71]. On the other hand, the resolution in the terahertz regime, corresponding to the gap scale of most noncentrosymmetric superconductors, [72] has recently been reported to reach 20 µrad and is expected to go down to 2 µrad. The optical rotation of TBG is close to this technical observable limit. Thus, unless we can find other two-dimensional materials with a larger Edelstein coefficient $\alpha$, the experimental observation is currently challenging in this regime.
@@
Finally, we shortly discuss the three-dimensional materials. In three-dimensional systems, the optical rotation caused by light transmitted through materials is expected. As light cannot easily penetrate superconductors, transmission can only occur in thin films [73]. This effect originates from the difference of the refractive index for left and right circular polarized lights $\Delta n$. It is given by $\Phi(=\theta+i\varepsilon)=-\Delta nk\zeta/2$, where $\theta$ and $\varepsilon$ are the rotational angle and the ellipticity, $k$ is the wavenumber of the light, and $\zeta$ is the thickness of materials. For the natural optical activity, the refractive index difference is given by $\Delta n=\mu_{0}c\chi$ [74]; thus, we can directly obtain the spectrum of the optical activity $\chi$ from the optical rotation spectrum. In the case of noncentrosymmetric cubics (point group: O and T), the optical activity in the THz regime is $\chi\sim(e\mu_{\mathrm{B}}\tau k_{F}^{2}/\hbar)(\gamma k_{F}/E_{F})$ with an antisymmetric SOC term $\gamma\bm{k}\cdot\bm{\sigma}$. Here, $\mu_{\mathrm{B}}$ is the Bohr magneton, $\tau$ is the relaxation time, and $k_{F}$ and $E_{F}$ are the Fermi wavenumber and the Fermi energy. Candidates for superconductors with this SOC include $\mathrm{Li_{2}Pd_{3}B}$, $\mathrm{Li_{2}Pt_{3}B}$, and $\mathrm{Mo_{2}Al_{3}C}$ [75]. Then, the optical rotation reads
$\Phi\sim-\frac{1}{2}\frac{e\mu_{\mathrm{B}}\mu_{0}k_{F}}{\hbar}(\Omega\tau)(k_{F}\zeta)\frac{\gamma k_{F}}{E_{F}},$ (23)
@@
and the typical order of magnitude is estimated to be 4.8 µrad for $k_{F}=1~{}\mathrm{\AA}^{-1}$, $\tau=10~{}\mathrm{fs}$, and $\gamma k_{F}/E_{F}=10^{-2}$ at $\Omega=1~{}\mathrm{THz}$ using the thickness $\zeta=500~{}\mathrm{\AA}$. It indicates that this optical rotation is observable in the THz regime [72].
@@ Appendix A No-go theorem
This equation is purely imaginary and antisymmetric under the interchange between $\mu$ and $\nu$. There is one difference to the case of the normal state. The velocity operator appearing by differentiating the Green’s function $\hat{v}^{\mu}_{\bm{k}}$ is different from the velocity operator $\tilde{v}^{\mu}_{\bm{k}}$. Thus, many components of the singular part can be finite.
Appendix B Optical rotation
@@
Here, we ignore the square of $\chi$ due to its smallness. This ratio gives the rotational angle $\theta$ and ellipticity $\varepsilon$ as
@@
The rotational angle and ellipticity of the transmitted light are obtained by simply replacing $r$ with $t$.